OpenStudy (anonymous):

Find the equation of the line that passes through the pair of points. Write your answer in slope-intercept form: (-2, 8), (1, -1)

OpenStudy (anonymous):

@whpalmer4 @mathmale @mathslover @ganeshie8

OpenStudy (anonymous):

Is slope-intercept form: y=mx+b ?

ganeshie8 (ganeshie8):

yes m = slope b = y intercept

ganeshie8 (ganeshie8):

start by finding the slope m

OpenStudy (anonymous):

\[m=-3\]

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@Kainui

OpenStudy (campbell_st):

well I think you need to re-work the slope calculations.

OpenStudy (anonymous):

\[m=\frac{ (y _{2}-y _{1}) }{ (x _{2}-x _{1}) }\] Right?

OpenStudy (campbell_st):

oops... sorry the slope is correct so the line is y = -3x + b substitute either point to find the value of b

OpenStudy (anonymous):

likw -2=-3(8)+b or just x or y by themselves?

OpenStudy (anonymous):

b=-22? @Hero

OpenStudy (anonymous):

\[\huge y-y _{1}=m(x-x _{1})\] You can use this and transform it into y =mx+b where y1 and x1 are (-2,8).

OpenStudy (anonymous):

y-8=m(x-(-2))?

OpenStudy (anonymous):

-8y=-3x-22? I'm lost

OpenStudy (anonymous):

@whpalmer4 @Hero @ganeshie8

OpenStudy (anonymous):

Sorry...OS doesn't work 90% of the time for me.

OpenStudy (anonymous):